OpenStudy (anonymous):

factor 56b2+23b-63 please so in detail how to factor.

OpenStudy (anonymous):

you need to use quadratic formula for this , because the numbers are far too big to factor it normally

OpenStudy (anonymous):

\[x =\frac{ -b +- \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

where a= 56, b=23 and c=-63

OpenStudy (anonymous):

after all that, you get solutions b=-9/7 and b= 7/8 therefore the factorisation is (b - (-9/7) )(b-(7/8))

OpenStudy (anonymous):

= ( 7b+9)(8b-7) after getting over common denominator etc